Grade 12Algebra

Solve the following:
The people of Sundargram planted trees in the village garden. Some of the trees were fruit trees. The number of non-fruit trees was two more than three times the number of fruit trees. What was the number of fruit trees planted if the number of non-fruit trees planted was 77?

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let the number of fruit tress be f.
From the question it is given that,
3 × number of fruit trees + 2 = number of non-fruit trees
= 3f + 2 = 77
By transposing 2 from LHS to RHS it becomes -2
=3f = 77–2
= 3f = 75
Divide both the side by 3
= 3f/3 = 75/3
= f = 25

Thanks